# EmoAttack: Emotion-to-Image Diffusion Models for Emotional Backdoor Generation

### v1

@@ 5.3 Comparison with Baselines
Table 3: Comparison of EmoBooth with Zero-day. Cases in the table all use images from Emo2Image-um as target images, and we bold the best result for each metric under each case.
@@
Sad Angry Isolated $Clip_{img1}^{tri}$uparrow $Clip_{img2}^{tri}$uparrow $Clip_{img3}^{tri}$uparrow Case1 EmoBooth 0.8360±0.0844 0.8335±0.0781 0.8162±0.0860 Zero-day 0.6882±0.0787 0.6905±0.0633 0.6509±0.0708 Case2 EmoBooth 0.7299±0.0788 0.7724±0.0719 0.6946±0.0635 Zero-day 0.6410±0.0738 0.6554±0.0742 0.6005±0.0534 Case3 EmoBooth 0.8121±0.0636 0.8083±0.0549 0.7808±0.0549 Zero-day 0.7654±0.0836 0.7798±0.0695 0.7589±0.0948 Case4 EmoBooth 0.8320±0.0758 0.8723±0.0449 0.8394±0.0612 Zero-day 0.6267±0.0970 0.6326±0.1193 0.5605±0.0450 Case5 EmoBooth 0.8041±0.0625 0.8506±0.0636 0.8024±0.0712 Zero-day 0.6868±0.0704 0.7232±0.0718 0.6534±0.0536
Table 4: Configured the same as in Table 3, except all selected images used as target images are from Emo2Image-m.
Comparison with Zero-day. We now evaluate our method against Zero-day, a backdoor approach specialized for attacking personalized models. As depicted in Tables 3 and 4, even after making some minor adjustments to Zero-day to better align with our task, the generated images under the backdoor scenario exhibit notable dissimilarity to the target images, resulting in significantly inferior outcomes compared to our approach. This distinction is further evident in the visual results presented in Fig. 5.
Statistical Analysis. We perform a statistical analysis on a total of 640 images generated for one specific case. As depicted in Fig. 4, in comparison to Censorship and Zero-day, the images generated by EmoBooth are closer to normal images under regular conditions and closer to target images under the backdoor scenario.
5.4 Visualization Comparison
Fig. 5 presents our visualization results, illustrating three emotions across four cases. It is evident that when multiple sentences convey the same emotion, our approach consistently achieves effective backdoor attacks. Additionally, when the provided target image corresponds to an object described in the text, our approach preserves good editability in the backdoor scenario. Furthermore, in situations where the text does not explicitly contain negative words but may imply negative emotions, our approach exhibits a certain level of recognition (e.g., “attack others"), indicating a more comprehensive ability in emotion recognition.
5.5 Discussion and Analysis
Effects of the number of texts for clustering. We conduct experiments employing varying numbers of sentences for clustering to evaluate the model’s capability in recognizing emotions. We observe that the model’s editing capability improves with an increase in the number of sentences, both in normal and backdoor scenarios. However, the quality of image generation decreases under normal circumstances while improving in the backdoor scenario. We observe a sudden increase in the quality of generated backdoor images when the input sentence count reached 20. This phenomenon is attributed to the optimal clustering of the 20 sentences, enhancing the identification of emotional centers and introducing a certain degree of randomness. Fig. LABEL:fig:4 (a) illustrates our analysis, which is displayed in the appendix.
Effects of the number of emotions. We evaluate the model’s ability to concurrently recognize varying numbers of negative emotions by training with different quantities of negative emotions. As depicted in  Fig. LABEL:fig:4 (b), we observe better performance in both editing capability and backdoor image generation when the number of emotion categories was set to 2. Conversely, under normal conditions, image quality decrease. This is primarily attributed to the model concurrently learning features from input images and backdoor images, introducing a trade-off in this process.
Probability value. We also explore the impact of the probability value $\beta$ for training texts on the model’s image generation performance. In  Fig. LABEL:fig:4 (c), with an increase in the probability value, the influence of normal images on the model parameters intensifies, leading to generated images that closely resemble normal images and deviate from the target image. When the probability value approaches 0.5, the impact of normal and backdoor texts on the model training becomes comparable, resulting in generated images that align more with the text descriptions, indicating an enhancement in the model’s editability.
6 Conclusion
In this work, we identified a new backdoor attack, i.e., EmoAttack, connecting the diffusion models with human motion, an essential element of the human experience. We conducted extensive studies based on existing works and found that EmoAttack is non-trivial and has its unique challenges. To tackle the challenges, we proposed a novel personalization method, i.e., EmoBooth, which incorporates emotion representation and emotion injection, allowing the targeted diffusion model to generate negative contents if specific emotion texts appear otherwise, producing normal images. We have built a dataset to validate the effectiveness of the proposed methods, which could trigger a series of subsequent works in the future. The results demonstrated the our method can properly achieve the EmoAttack and outperform baselines significantly.
7 Impact Statements
@@
Our research is focused on uncovering vulnerabilities associated with exploiting user emotions as a backdoor, resulting in the generation of malicious specified images by diffusion models. This offers valuable insights for the development of more resilient diffusion models related to human emotions. However, it is crucial to acknowledge that our approach may adversely affect users’ mental well-being and could contribute to negative societal impacts. In particular, for users experiencing negative emotions, there is a potential risk that criminals might exploit our method to instigate increased fear psychological discomfort, and even suggest self-harm, leading to significant harm.

### v3

@@ D.5 Generalization to Other Emotion Types
TABLE VI: Comparison with Censorship under the metrics of Clip Score and EmoAttack Capability (EAC). Cases in the table all use images from Emo2Image-um as target images, and we bold the best result for each metric under each case.
EAC $\uparrow$ Annoyed Nervous Scared Normal $\text{Clip}_{\text{txt1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt}}$ $\uparrow$ $\text{Clip}_{\text{img}}$ $\uparrow$ Set1 EmoBooth 0.6539 0.2587±0.0257 0.8325±0.0633 0.2457±0.0236 0.8420±0.0787 0.2533±0.0170 0.8529±0.0697 0.2538±0.0385 0.7230±0.0639 Censorship 0.6182 0.2380±0.0190 0.7823±0.0835 0.2328±0.0199 0.8025±0.0734 0.2388±0.0304 0.7923±0.0846 0.2419±0.0363 0.7130±0.0737 Set2 EmoBooth 0.6369 0.2653±0.0325 0.8350±0.0525 0.2532±0.0221 0.8128±0.0797 0.2485±0.0326 0.8016±0.0508 0.2571±0.0215 0.7015±0.0639 Censorship 0.5981 0.2358±0.0295 0.7725±0.0549 0.2266±0.0388 0.7358±0.0839 0.2462±0.0222 0.7528±0.0860 0.2642±0.0176 0.7225±0.0532 Set3 EmoBooth 0.6206 0.2538±0.0236 0.7726±0.0838 0.2389±0.0370 0.7820±0.0747 0.2587±0.0331 0.7923±0.0821 0.2566±0.0177 0.7552±0.0846 Censorship 0.5540 0.2650±0.0231 0.6859±0.0607 0.2358±0.0341 0.6849±0.0760 0.2318±0.0317 0.6523±0.0609 0.2533±0.0313 0.7520±0.0899 Set4 EmoBooth 0.6435 0.2432±0.0344 0.8624±0.0807 0.2532±0.0341 0.8532±0.0827 0.2311±0.0230 0.8458±0.0543 0.2358±0.0197 0.5918±0.0723 Censorship 0.5891 0.2380±0.0312 0.7599±0.0860 0.2158±0.0331 0.7836±0.0732 0.2189±0.0373 0.7520±0.0884 0.2312±0.0295 0.6213±0.0509 Set5 EmoBooth 0.6620 0.2610±0.0297 0.8720±0.0554 0.2432±0.0260 0.8521±0.0877 0.2321±0.0182 0.8629±0.0657 0.2519±0.0179 0.7042±0.0592 Censorship 0.5988 0.2258±0.0198 0.7856±0.0540 0.2385±0.0174 0.7325±0.0545 0.2178±0.0336 0.7628±0.0848 0.2699±0.0295 0.7019±0.0791
TABLE VII: Configured as in Table VI, except for the Sets in the table using cases from Emo2Image-m as target images, the weighting coefficient for EAC is different, and here, we aim for higher values in $\text{Clip}_{\text{txt}}^{\text{tri}}$.
D.6 Influence of $\lambda$ in Eq. (8)
In Eq. (8), We set $\lambda=1$ primarily to balance the weights between prior knowledge and input image features. Here, we conducted ablation experiments by evaluating the CLIP score under different $\lambda$ values in both normal and backdoor scenarios. As shown in Fig. 18, when $\lambda<1$, the CLIP scores for both normal and backdoor scenarios are relatively low, especially the CLIP text score. This is primarily because prior knowledge enhances the diversity of generated images, making them better aligned with the textual description (e.g., generating various poses of a dog). However, when $\lambda>1$, the CLIP image score decreases rapidly. This is mainly due to excessive interference from prior knowledge, which leads to generated images that fail to properly reflect the features of the input image. Therefore, we chose $\lambda=1$ as the balance point.
Figure 18: Influence of $\lambda$ in Eq. (8).
D.7 Results against the latest stable diffusion model
EmoBooth was originally implemented using Stable Diffusion v1.4. We have reconstructed EmoBooth based on Stable Diffusion v2.1 and conducted experiments under the EmoSet-m scenario. As shown in Table VIII, EmoBooth achieves the highest EAC score compared to the baseline, even with the v2.1 version of Stable Diffusion. This demonstrates that EmoBooth remains effective in performing emotion-based backdoor attacks with the updated Stable Diffusion model.
@@ D.7 Results against the latest stable diffusion model
EAC $\uparrow$ Sad Angry Isolated Normal $\text{Clip}_{\text{txt1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt}}$ $\uparrow$ $\text{Clip}_{\text{img}}$ $\uparrow$ Set1 EmoBooth 0.6511 0.2690±0.0317 0.8360±0.0844 0.2532±0.0421 0.8521±0.0538 0.2510±0.0251 0.8232±0.0738 0.2585±0.0214 0.7150±0.0482 Censorship 0.6060 0.2870±0.0318 0.7822±0.0884 0.2331±0.0244 0.7705±0.0691 0.2497±0.0251 0.7431±0.0892 0.2428±0.0292 0.7130±0.0588 Set2 EmoBooth 0.5894 0.2420±0.0332 0.7421±0.0788 0.2382±0.0165 0.7638±0.0719 0.2577±0.0318 0.7214±0.0642 0.2574±0.0302 0.6910±0.0900 Censorship 0.5666 0.2453±0.0333 0.6776±0.0589 0.2463±0.0209 0.7362±0.0678 0.2406±0.0284 0.6758±0.0515 0.2616±0.0298 0.7373±0.0694 Set3 EmoBooth 0.6396 0.2655±0.0257 0.8232±0.0732 0.2438±0.0212 0.8128±0.0549 0.2543±0.0258 0.8023±0.0430 0.2477±0.0280 0.7628±0.0653 Censorship 0.6270 0.2580±0.0296 0.7966±0.0532 0.2624±0.0196 0.8075±0.0494 0.2529±0.0339 0.7682±0.0646 0.2509±0.0329 0.7558±0.0649 Set4 EmoBooth 0.6372 0.2543±0.0542 0.8732±0.0677 0.2343±0.0125 0.8280±0.0538 0.2428±0.0386 0.8366±0.0712 0.2318±0.0187 0.5777±0.0613 Censorship 0.5936 0.2108±0.0357 0.7422±0.0564 0.2169±0.0206 0.8165±0.0548 0.2248±0.0303 0.7392±0.0697 0.2198±0.0329 0.6851±0.0329 Set5 EmoBooth 0.6491 0.2534±0.0432 0.8353±0.0628 0.2370±0.0312 0.8706±0.0572 0.2428±0.0251 0.8143±0.0712 0.2433±0.0286 0.7188±0.0709 Censorship 0.6332 0.2480±0.0362 0.7908±0.0626 0.2428±0.0203 0.8602±0.0420 0.2605±0.0247 0.7809±0.0523 0.2638±0.0285 0.7040±0.0587
TABLE VIII: Using Stable Diffusion v2.1, we constructed EmoBooth, with all experimental datasets sourced from EmoSet-m.
Appendix E More Discussions for EmoBooth
We discussed broader and potentially malicious applications of EmoBooth, and its achievable positive impacts.
EAC $\uparrow$ Happy Optimistic Enthusiastic Normal $\text{Clip}_{\text{txt1}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt2}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt3}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt}}$ $\uparrow$ $\text{Clip}_{\text{img}}$ $\uparrow$ Set1 EmoBooth 0.7715 0.1897±0.0437 0.7421±0.0923 0.2407±0.0418 0.7865±0.1240 0.2534±0.0372 0.7652±0.1157 0.2562±0.0292 0.7708±0.0756 Censorship 0.7326 0.2631±0.0350 0.7146±0.0411 0.2477±0.0320 0.7291±0.0902 0.2544±0.0306 0.7226±0.0837 0.2521±0.0271 0.7774±0.0709 Set2 EmoBooth 0.7296 0.1708±0.0584 0.7365±0.0788 0.2378±0.0507 0.7472±0.0719 0.2296±0.0504 0.6475±0.0635 0.2546±0.0287 0.7646±0.0997 Censorship 0.6118 0.1836±0.0520 0.5967±0.1161 0.2559±0.0370 0.5604±0.0812 0.2697±0.0275 0.5489±0.0547 0.2605±0.0296 0.7603±0.0648 Set3 EmoBooth 0.8210 0.1707±0.0513 0.8392±0.1126 0.1403±0.0426 0.8459±0.0839 0.1583±0.0426 0.8364±0.0999 0.2284±0.0414 0.6710±0.1111 Censorship 0.6374 0.2554±0.0432 0.6418±0.1221 0.2375±0.0380 0.6251±0.0918 0.2503±0.0384 0.6059±0.0989 0.2569±0.0288 0.6810±0.1050 Set4 EmoBooth 0.8474 0.1120±0.0518 0.8761±0.1125 0.1043±0.0382 0.8899±0.0931 0.1291±0.0629 0.8384±0.1714 0.1980±0.0758 0.6816±0.2086 Censorship 0.6704 0.1591±0.0904 0.7510±0.2389 0.1971±0.0826 0.6108±0.2578 0.2030±0.0767 0.5728±0.2502 0.2394±0.0576 0.7194±0.1623 Set5 EmoBooth 0.8118 0.2376±0.0511 0.7908±0.1450 0.2186±0.0498 0.8602±0.1520 0.2382±0.0364 0.7809±0.1325 0.2494±0.0314 0.7985±0.0737 Censorship 0.6460 0.2537±0.0504 0.6125±0.1188 0.2424±0.0363 0.6294±0.0806 0.2473±0.0397 0.5980±0.1391 0.2572±0.0336 0.7672±0.0829
TABLE IX: Comparison with Censorship using positive emotions as trigger. Sets in the table all use cases from Emo2Image-um as target images, and we bold the best result for each metric under each Set.
EAC $\uparrow$ Happy Optimistic Enthusiastic Normal $\text{Clip}_{\text{txt1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{img3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt}}$ $\uparrow$ $\text{Clip}_{\text{img}}$ $\uparrow$ Set1 EmoBooth 0.6097 0.2642±0.0385 0.7948±0.0608 0.2528±0.0216 0.7636±0.0561 0.2622±0.0227 0.7357±0.0508 0.2499±0.0270 0.7401±0.0777 Censorship 0.5911 0.2488±0.0358 0.7548±0.1740 0.2427±0.0207 0.7345±0.1980 0.2534±0.0219 0.7278±0.1648 0.2520±0.0278 0.7277±0.0747 Set2 EmoBooth 0.5748 0.2562±0.0330 0.7188±0.0694 0.2543±0.0228 0.7016±0.0524 0.2578±0.0246 0.6964±0.0545 0.2489±0.0384 0.7534±0.0975 Censorship 0.5687 0.2604±0.0529 0.7155±0.1460 0.2450±0.0414 0.6990±0.1340 0.2565±0.0379 0.6774±0.1422 0.2532±0.0292 0.7421±0.0752 Set3 EmoBooth 0.6115 0.2579±0.0339 0.7724±0.0424 0.2522±0.0252 0.7820±0.0366 0.2632±0.0193 0.7678±0.0289 0.2318±0.0537 0.7231±0.0923 Censorship 0.6035 0.2547±0.0553 0.7489±0.1046 0.2425±0.0539 0.7599±0.1382 0.2479±0.0466 0.7656±0.1515 0.2531±0.0535 0.7366±0.1259 Set4 EmoBooth 0.6351 0.2541±0.0355 0.8375±0.0361 0.2125±0.0277 0.8424±0.0434 0.2189±0.0275 0.8377±0.0407 0.2147±0.0344 0.6443±0.0738 Censorship 0.6206 0.2404±0.0338 0.8048±0.0535 0.2300±0.0266 0.8223±0.0520 0.2414±0.0238 0.8036±0.0616 0.2354±0.0322 0.6344±0.0779 Set5 EmoBooth 0.6606 0.2343±0.0303 0.8577±0.0242 0.2586±0.0242 0.8651±0.0636 0.2388±0.0217 0.8610±0.0712 0.2411±0.0264 0.7099±0.0634 Censorship 0.6353 0.2667±0.0338 0.8229±0.0521 0.2539±0.0223 0.8230±0.0527 0.2688±0.0192 0.7956±0.0559 0.2378±0.0290 0.7057±0.0611
TABLE X: Configured as in Table IX, except for the Sets in the table using cases from Emo2Image-m as target images, the weighting coefficient for EAC is different, and here, we aim for higher values in $\text{Clip}_{\text{txt}}^{\text{tri}}$.
The inference of using positive emotions. In addition to negative emotions, we also employed positive emotions as triggers for comparative experiments to showcase EmoBooth’s effectiveness in targeting a variety of emotions. To provide a comprehensive assessment, we selected three emotions: happiness, optimism, and enthusiasm, and conducted experiments accordingly. The results, as depicted in Tables IX and X show that, akin to using negative emotions as triggers, our method achieved optimal effectiveness.
Figure 19: Visualization results using negative contents for specific populations.
Targeted attacks on specific demographics. Here, we showcase potential malicious applications of our attacks. For instance, attackers could initially profile users and categorize them based on their backgrounds, enabling targeted malicious assaults. Fig. 19 illustrates four specific user profiles and the corresponding generated outcomes, including bloody phobia, soldier, student, and depression patients. Target contents are set as bloody images, war images, bullying images, and suicide suggestive images, respectively, to showcase the malicious applications inflicting psychological trauma on users.
Figure 20: Visualization results using positive contents for specific populations.
Positive influence of our Work. We also explore the positive applications of our method. Indeed, we can readily replace targeted negative contents with positive ones. This approach allows us to associate specific emotions, such as negative ones, with targeted positive content. Fig. 20 demonstrates the therapeutic effects of our work on the minds of specific demographics. We selected four groups: individuals experiencing depression, soldiers, lonely individuals, and children with autism. We replaced the target contents with images beneficial to the psychological well-being of these groups to showcase the positive applications of our work.
Attack effectiveness varies with input cases. Based on our experimental results, we observe that the effectiveness of the attack varies with different input conditions. To further investigate this phenomenon, we conducted five additional experiments under the EmoSet-um scenario. As illustrated in Table XI, when the input image is from set1, the CLIP text scores for the three emotional prompts fluctuate around 0.19. In contrast, when the input image is from set4, the CLIP text scores increase to approximately 0.24. Similarly, the CLIP image scores fluctuate around 0.73 for images from set4 but rise to approximately 0.83 for images from set5.
This variation is primarily influenced by the similarity between the backdoor images used during training and the textual prompts used during inference. Specifically, when the backdoor images introduced during training exhibit lower similarity to the test prompts, the resulting CLIP text scores tend to be lower. Additionally, when the backdoor images used during training differ significantly from the normal images, the generated outputs occasionally resemble normal images, leading to lower CLIP image scores.
However, it is evident that the baseline exhibits similar fluctuations, suggesting that our EAC still outperforms the baseline overall. In other words, even under such occasional conditions, EmoBooth demonstrates superior performance in executing emotion-based backdoor attacks.
EAC $\uparrow$ Sad Angry Isolated Normal $\text{Clip}_{\text{txt1}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt2}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt3}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt}}$ $\uparrow$ $\text{Clip}_{\text{img}}$ $\uparrow$ Set1 EmoBooth 0.7124 0.1928±0.0313 0.7928±0.1231 0.2058±0.0425 0.8635±0.1248 0.1932±0.0230 0.8488±0.1644 0.2532±0.0468 0.586±0.1377 Censorship 0.6242 0.2233±0.0183 0.6739±0.1853 0.2336±0.0261 0.7023±0.1738 0.2015±0.0249 0.7856±0.1857 0.2641±0.0313 0.6518±0.0955 Set2 EmoBooth 0.7059 0.1843±0.0277 0.7963±0.1533 0.1857±0.0265 0.8429±0.1328 0.1732±0.0347 0.8133±0.1623 0.2081±0.0275 0.6732±0.1414 Censorship 0.5758 0.2024±0.0277 0.6243±0.1228 0.2143±0.0238 0.6165±0.1421 0.2236±0.0311 0.7358±0.1177 0.2205±0.0287 0.6923±0.0923 Set3 EmoBooth 0.7147 0.1963±0.0128 0.8082±0.0938 0.1976±0.0211 0.8617±0.0788 0.1829±0.0253 0.8023±0.1142 0.2370±0.0533 0.7021±0.1251 Censorship 0.5922 0.2141±0.0229 0.6758±0.1281 0.2242±0.0231 0.7218±0.1532 0.2423±0.0377 0.6759±0.1120 0.1937±0.0326 0.6535±0.1231 Set4 EmoBooth 0.6392 0.2356±0.0432 0.7324±0.1827 0.2436±0.0228 0.7336±0.1129 0.2421±0.0319 0.7523±0.1539 0.2343±0.0283 0.7236±0.1872 Censorship 0.5754 0.2453±0.0298 0.6828±0.1927 0.2587±0.0312 0.6658±0.1765 0.2578±0.0283 0.657±0.1927 0.2217±0.0476 0.6923±0.1326 Set5 EmoBooth 0.7309 0.1984±0.0432 0.8644±0.1687 0.1950±0.0287 0.8351±0.1333 0.1925±0.0425 0.8267±0.1187 0.2458±0.0287 0.7168±0.1277 Censorship 0.5995 0.2242±0.0381 0.6728±0.1577 0.2381±0.0276 0.7123±0.1382 0.1966±0.0299 0.6925±0.1281 0.2316±0.0370 0.6623±0.0841
TABLE XI: Evaluating the Impact of Input Images on Experimental Results: All Experimental Datasets Are Derived from EmoSet-um.
Appendix F Safety and Ethical Statement
@@
The EmoBooth project adheres to strict safety and ethical standards throughout the development, deployment, and dissemination of its Emotion-Based Backdoor Attack Propagation Model and EmoSet dataset. Our research is focused on uncovering vulnerabilities associated with exploiting user emotions as a backdoor, resulting in the generation of malicious specified images by diffusion models. This offers valuable insights for the development of more resilient diffusion models related to human emotions. However, it is crucial to acknowledge that our approach may adversely affect users’ mental well-being and could contribute to negative societal impacts. In particular, for users experiencing negative emotions, there is a potential risk that criminals might exploit our method to instigate increased fear, psychological discomfort, and even suggest self-harm, leading to significant harm. The following points outline the measures and considerations taken to ensure the responsible and ethical use of our work:
  1. 1.
    Targeted Vulnerable Models: Our attack model is specifically designed to demonstrate vulnerabilities in text-to-image diffusion models such as Stable Diffusion, ControlNet, and Glide. It is intended for research, educational, and lawful security testing purposes. We unequivocally condemn any attempt to employ our attack methods for malicious or unauthorized activities.
  2. 2.
    Controlled Release of Code and Dataset: To ensure that our code and dataset are accessed and used responsibly, we have implemented a rigorous controlled release mechanism:
    1. (a)
      Application-Based Access: Access to the EmoSet dataset and code will be granted only through a formal application process. Interested researchers must submit a detailed application explaining their intended use, research objectives, and the security measures they will implement.
    2. (b)
      Review and Approval: A dedicated review committee will evaluate each application based on strict ethical standards, security protocols, and potential societal impact. Access will only be granted to legitimate research institutions and verified researchers who demonstrate a strong commitment to ethical practices.
    3. (c)
      Regular Audits: Researchers granted access will be subject to periodic audits to ensure adherence to agreed-upon terms and conditions. Any breach of compliance may result in revocation of access and potential legal actions.
  3. 3.
    User Agreement and Responsibility: Researchers seeking access to the EmoSet dataset and model code must agree to the following conditions:
    1. (a)
      Signing a Legally Binding Agreement: Prior to access, researchers will sign a legal document outlining the terms of use, which includes restrictions on data sharing, obligations to report any security breaches, and adherence to ethical guidelines.
    2. (b)
      Commitment to Ethical Conduct: Users must commit to conducting their research in accordance with the highest standards of ethics, ensuring respect for privacy, and avoiding any action that could harm individuals or groups.
    3. (c)
      Liability Clause: The agreement includes a liability clause, making researchers accountable for any misuse or unauthorized dissemination of the dataset or code.
  4. 4.
    Secure Distribution and Monitoring: To maintain the security and integrity of the EmoSet dataset and code, we employ the following measures:
    1. (a)
      Secure Distribution Channels: All data and code are distributed through encrypted channels, requiring multi-factor authentication to ensure that only approved researchers can access the materials.
    2. (b)
      Access Tracking: A sophisticated access tracking system monitors all usage of the dataset and code. Detailed logs, including access timestamps and user identities, are maintained to prevent unauthorized access and ensure accountability.
    3. (c)
      Regular Usage Reports: Researchers are required to submit regular reports detailing their use of the dataset and code. These reports will be reviewed by the committee to ensure compliance with the terms of access.
  5. 5.
    Ethical Data Collection: The images in EmoSet were sourced following ethical guidelines and strict copyright considerations:
    1. (a)
      Data Sources: Images were collected from three websites (Baidu, Playground, Yandex) as detailed in Appendix B.2. Images were manually curated, and any human-related content generated using diffusion models without safety checks was reviewed to ensure ethical standards.
    2. (b)
      Copyright Compliance: We have reviewed the terms of use for the images from these sources:
      1. i.
        Images from Playground were used in accordance with their open creative community policy.
      2. ii.
        Images from Yandex and Baidu were used with strict adherence to non-commercial terms.
      3. iii.
        Any third-party web-linked images underwent a copyright verification process.
  6. 6.
    Reporting and Mitigating Vulnerabilities: We encourage all users to report any discovered vulnerabilities or issues related to EmoBooth promptly. Users must cooperate fully with investigations to resolve these issues and help prevent potential misuse.